# Covariate-Adaptive Randomization in Clinical Trials Without Inflated Variances

### v1

@@ Abstract
Covariate adaptive randomization (CAR) procedures are extensively used to reduce the likelihood of covariate imbalances occurring in clinical trials. In literatures, a lot of CAR procedures have been proposed so that the specified covariates are balanced well between treatments. However, the variance of the imbalance of the unspecified covariates may be inflated comparing to the one under the simple randomization. The inflation of the variance causes the usual test of treatment effects being not valid and adjusting the test being not an easy work. In this paper, we propose a new kind covariate adaptive randomization procedures to balance covariates between two treatments with a ratio $\rho:(1-\rho)$. Under this kind of CAR procedures, the convergence rate of the imbalance of the specified covariates is $o(n^{1/2})$, and at the same time the asymptotic variance of the imbalance of any unspecified (observed or unobserved) covariates does not exceed the one under the simple randomization. The ”shift problem” found by Liu, Hu, and Ma (2025) will not appear under the new CAR procedures.
@@ 1 Introduction
Covariate adaptive randomization (CAR) procedures have often been implemented for balancing treatment allocation over influential covariates in clinical trials (McEntegart, , 2003; Rosenberger and Lachin, , 2015; Taves, , 2010; Lin, Zhu and Su, , 2015). The main purposes are to enhance the comparability of treatment groups and to increase the efficiency of the statistical inference. Many CAR procedures have been proposed to balance treatment assignments within strata and over covariate margins, provided that the covariates under consideration are discrete (categorical with two or more levels) (Pocock and Simon, , 1975; Antognini and Zagoraiou, , 2011; Hu and Hu, , 2012; Rosenberger and Lachin, , 2015). The theories on CAR randomization procedures are established progressively in the past decade(Hu and Hu, , 2012; Ma, Hu and Zhang, , 2015; Hu and Zhang, , 2020; Hu, Ye and Zhang, , 2023). For balancing continuous covariates, besides the method of discretization, a variety of methods have been proposed by minimization of prescribed imbalance measures (Hoehler, , 1987; Nishi and Takaichi, , 2004; Endo et al., , 2006; Stigsby and Taves, , 2010; Ciolino et al., , 2011; Lin and Su, , 2012; Ma and Hu, , 2013; Qin et al., , 2018; Jiang, Ma and Yin, , 2018; Li, Zhou and Hu, , 2019). Recently, Ma et al, (2024) unified a lot of CAR procedures by proposing a general CAR procedure for balancing specified covariate features $\bm{\phi}(\bm{X}_{i})$ between two treatments $1$ and $2$, where $\bm{\phi}(\bm{X}_{i})=(\phi_{1}(\bm{X}_{i}),\ldots,\phi_{q}(\bm{X}_{i}))$ can contain additional covariate features besides the covariate vector $\bm{X}_{i}=(X_{i,1},\ldots,X_{i,p})$ of a unit in clinical trals, such as quadratic and interaction terms. It is shown that the convergence rate of the imbalance vectors $\sum_{i=1}^{n}(2T_{i}-1)\bm{\phi}(\bm{X}_{i})$ is $o_{P}(n^{\epsilon_{0}})$, $\epsilon_{0}<1/2$, where $n$ is the sample size, $\bm{X}_{i}$ is the covariate vector of the $i$-th unit, and $T_{i}=1$, 0 stands for the $i$-unit being allocated to treatment $1$ and $2$, respectively. Under some suitable conditions on the density of $\bm{\phi}(\bm{X}_{i})$, $\sum_{i=1}^{n}(2T_{i}-1)\bm{\phi}(\bm{X}_{i})$ can attain the fastest convergence rate $O_{P}(1)$, and at the same time, for any unspecified covariate feature $m(\bm{X}_{i})$, the asymptotic normality of its normalized imbalance $\sum_{i=1}^{n}(2T_{i}-1)m(\bm{X}_{i})/\sqrt{n}$ holds and the asymptotic variance $\sigma_{m}^{2}$ is nonzero when $m(\bm{X}_{i})$ is not a linear function of $\bm{\phi}(\bm{X}_{i})$. However, $\sigma_{m}^{2}$ may exceed $\textsf{E}[m^{2}(\bm{X}_{i})]$ that is the variance under the complete randomization in which each unit is allocated to treatments $1$ and $2$ with equal probability $1/2$. This issue is referred to as the inflation of variance. The inflation of variance under the Pocock and Simon, (1975) is also founded by Zhao et al., (2024). Because the asymptotic variance of the test of treatment effects is usually a function of $\sigma_{m}^{2}$ for some unspecified covariate feature $m(\bm{X}_{i})$, the inflation of variance causes such test being not valid (Ma et al, , 2024; Zhang, , 2023). The closed form of $\sigma_{m}^{2}$ is also unknown so that estimating it and adjusting the test are not easy works. A lot of methods for the adjustment to obtain robust statistical inferences in the CAR clinical trials have being proposed(Bugni, Canay and Shaikh, , 2018; Ye and Shao, , 2020; Liu and Yang, , 2020; Ma, Tu and Liu, , 2024; Wang et al., , 2023). The purpose of this paper is to propose a new kind of CAR procedures so that the specified covariate features $\bm{\phi}(\bm{X}_{i})$ can be balanced well and at the same time the asymptotic variance $\sigma_{m}^{2}$ of the normalized imbalance of any unspecified covariate feature $m(\bm{X}_{i})$ is not inflated and has a closed form.
@@
From the proofs (Hu and Zhang, , 2020; Ma et al, , 2024; Zhang, , 2023), we can find that the convergence of $\sum_{i=1}^{n}(2T_{i}-1)m(\bm{X}_{i})/\sqrt{n}$ and the finiteness of its asymptotic variance $\sigma_{m}^{2}$ are strongly based on the theory of Markov Chains and the property that the allocation probabilities are symmetric about treatments so that, as a Markov Chain, $\sum_{i=1}^{n}(2T_{i}-1)\bm{\phi}(\bm{X}_{i})$ has a symmetric invariant distribution and thus $\sum_{i=1}^{n}(2T_{i}-1)m(\bm{X}_{i})/\sqrt{n}$ has a zero mean under the invariance distribution for any $m(\bm{X}_{i})$. Recently, Liu, Hu and Ma, (2025) generalized Ma et al, (2024)’s procedure to balance covariances between two treatments with a ratio $\rho:(1-\rho)$ instead of half to half. The convergence rate of the imbalance vectors $\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})$ of the specified covariate feature $\bm{\phi}(\bm{X}_{i})$ is also shown to be $o_{P}(n^{\epsilon_{0}})$, $\epsilon_{0}<1/2$. However, by their theory, for an unspecified covariate feature $m(\bm{X})$, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})/n$ may converge to a nonzero constant $c_{m}$ due to the fact that $\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})$ is no longer a symmetric Markov Chain when $\rho\neq 1/2$. This issue is refer to as ”shift problem” of covariates $m(\bm{X}_{i})$. The shift problem is a bad news because in testing hypotheses in the covaraite-adaptive randomized clinical trials, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})=o_{P}(n)$ is a basic condition (Ma, Hu and Zhang, , 2015; Ma et al., , 2020; Ma et al, , 2024; Zhu et al, , 2025; Liu and Hu, , 2023; Liu, Hu and Ma, , 2025), and sometimes the convergence rate $O_{P}(\sqrt{n})$ is needed (Ma, Tu and Liu, , 2024). We will show that our proposed CAR procedures are valid for all the case of $0<\rho<1$. Under the proposed CAR procedure, the ”shift problem” will no longer appear and the asymptotic variance $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})/\sqrt{n}$ will be always controlled by $\textsf{E}[m^{2}(\bm{X}_{i})]$ for any $m(\bm{X}_{i})$.
@@ 2.1 Framework
    $\textsf{P}(T_{n}=1|\bm{X}_{n},...,\bm{X}_{1},T_{n-1},...,T_{1})=\ell\left(\frac{\langle\bm{\Lambda}_{n-1},\bm{\phi}(\bm{X}_{n})\rangle}{(n-1)^{\gamma}}\right),$ (2.1)
@@
The function $\ell(x)$ in (2.1) is called the allocation function. A simple choice of $\ell$ is
@@
where $x\wedge y=\min\{x,y\}$ and $x\vee y=\max\{x,y\}$.
According to the theoretical results given in Theorem 2.2, the parameter $\gamma$ controls the size of the imbalance measure $Imb_{n}$. The smaller the parameter $\gamma$ , the smaller is $Imb_{n}$, but possibly the larger is the imbalance $\|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\|^{2}$ of the unspecified feature $Z_{i}$ that is not a linear fuction of $\bm{\phi}(\bm{X}_{i})$. The value of $\gamma$ is suggested to be in $(0.5,1)$ so that the asymptotic expectation of $\|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\|^{2}/n$ does not exceed the one under the simple randomization for any $Z_{i}$s.
@@ 2.2 Theoretical Properties
The sequence $\{(\bm{X}_{i},Z_{i},W_{i})\}$ is a sequence of independent and identically distributed random vectors.
@@
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{2}]=O\big(n^{\gamma}+n^{\frac{1}{2}+\frac{1}{2(r_{0}-1)}}\big)=o(n),$ (2.5)
and
$\frac{\textsf{E}\left|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\right|}{n}\to 0$ (2.6)
if $\textsf{E}[|Z|]<\infty$ and $\gamma>\frac{1}{4}+\frac{1}{4(r_{0}-1)}$. Moreover, if $r_{0}\geq 4$, $\gamma>\frac{1}{2}+\frac{1}{4(r_{0}-1)}$, $\textsf{E}Z^{2}<\infty$, $\textsf{E}W^{2}<\infty$ and $\textsf{E}[W]=0$, then
@@
Liu, Hu and Ma, (2025) proposes a CAR procedure for balancing covariate features $\bm{\phi}(\bm{X}_{i})$ between treatments with an unequal ratio $\rho:(1-\rho)$. When the additional covariate $m(\bm{X}_{i})$ is not the linear function of $\bm{\phi}(\bm{X}_{i})$, it may happen that $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})\overset{P}{\to}c\neq 0$. This issue is referred as ”shift problem”. The simulation results presented in the Supplementary Material of Liu, Hu and Ma, (2025) show that the shift problem on the additional covariates is common when balancing continuous covariates by using their CAR procedure. Now, according to (2.6), the shift problem no longer exists under our proposed CAR procedure.
$\langle\bm{\phi}(\bm{X}),\bm{x}_{0}\rangle$ is the orthogonal projection of $Z$ on $\bm{\phi}(\bm{X})$, we denote it by $\textsf{Pj}[Z|\bm{\phi}(\bm{X})]$. It is easily seen that $\textsf{E}\big(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})]\big)^{2}=\textsf{E}Z^{2}-\textsf{E}\big(\textsf{Pj}[Z|\bm{\phi}(\bm{X})]\big)^{2}\leq\textsf{E}Z^{2}$. Thus, $\vec{\sigma}_{Z}^{2}=0$ if and only if $Z\in Span\{\bm{\phi}(\bm{X})\}=:\{\langle\bm{x},\bm{\phi}(\bm{X})\rangle:\bm{\in}\mathbb{R}^{q}\}$, and $\vec{\sigma}_{Z}^{2}=(1-\rho)\rho\textsf{E}[Z^{2}]$ if and only if $\textsf{E}[\bm{\phi}(\bm{X})Z]=\bm{0}$, i.e., $Z\perp\bm{\phi}(\bm{X})$. As a result, $\vec{\sigma}_{Z}^{2}=\textsf{Var}(Z-\textsf{E}[Z|\bm{\phi}(\bm{X})])+\vec{\sigma}_{\textsf{E}[Z|\bm{\phi}(\bm{X})]}^{2}$. Further, if $1\in Span\{\bm{\phi}(\bm{X})\}$, then $\vec{\sigma}_{Z}^{2}=\vec{\sigma}_{Z-\textsf{E}Z}\leq\textsf{Var}(Z)$.
The simple randomization is randomizing each patient to treatment $1$ with the same probability $\rho$. For the simple randomization, $\vec{\sigma}_{Z}^{2}=\textsf{E}Z^{2}$. So, the proposed CAR procedure always has smaller asymptotic variance than the simple randomization for all observed or unobserved feature $Z_{i}$ with $\textsf{Pj}[Z|\bm{\phi}(\bm{X})]\neq 0$. The inflation of variance is no longer a problem under our proposed CAR procedure.
(2.7) means that the total sum of $W_{i}$s is asymptotic independent of the imbalance of $Z_{i}$s regardless $W_{i}$ and $Z_{i}$ themselves are independent or not.
When $\bm{\phi}(\bm{X})$ has finite moments of all orders. The convergence rate of the imbalance $\Lambda_{n}$ can be improved a lot.
@@
Suppose Assumption 2.1, and that 2.2 is satisfied for all $r_{0}>2$. Then
$\big(\textsf{E}\|\bm{\Lambda}_{n}\|^{r}\big)^{1/r}=O(n^{\gamma/2})\;\text{ for all }r>0,$ (2.8)
(2.6) holds if $\textsf{E}[|Z|]<\infty$, and
@@
for all $r>0$, if $\textsf{E}Z^{2}<\infty$, where $\widetilde{Z}_{i}=Z_{i}-\textsf{Pj}[Z_{i}|\bm{\phi}(\bm{X}_{i})]$. Moreover, if $\textsf{E}Z^{2}<\infty$, $\textsf{E}W^{2}<\infty$, $\textsf{E}W=0$ and $1/2<\gamma<1$, then (2.7) holds,
@@
If $\ell(x)$ is thrice differentiable at $x=0$, and $\ell^{\prime\prime}(0)=0$ or $\textsf{E}[(\bm{\phi}(\bm{X}))^{\prime}\bm{\phi}(\bm{X})\widetilde{Z}]=\bm{0}$, then $n^{1-\gamma}$ in (2.2) can be repalced by $n^{1-3\gamma/2}$, and (2.7), (2.10), (2.11) remain true when $1/3<\gamma<1$.
@@
where $\lambda>0$ is a tuning parameters. In (2.12), $\Phi(-x+u_{\rho/2})$, $\Phi(-x+u_{(\rho+1)/2})$ can be replaced by $\alpha\rho$, $1-(1-\alpha)\rho$, respectively, $0<\alpha<1$.
@@
where $w_{o}$, $w_{m,l}$, $w_{s}$ are the corresponding nonnegative weights.
@@
$=$ $w_{0}D_{n}+\sum_{l=1}^{p}D_{n-1}(l;k_{l}^{\ast})+w_{s}D_{n-1}(k_{1}^{\ast},\ldots,k_{L}^{\ast}).$
Then, the allocation probability of the $n$-unit is
@@
Thus, when $w_{s}\neq 0$, $D_{n}^{2}(k_{1},\ldots,k_{L})=O_{P}(n^{\gamma/2})$ and thus for any $m(\bm{X})$, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})=O_{P}(n^{\gamma/2})$. When $w_{s}=0$, we will have
@@
when $X_{1},\ldots,X_{p}$ have finite moments of each order. At the same time, for any nonlinear function $m=m(X_{1},\ldots,X_{p})$ with finite variance, then the asymptotic variance of $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})/\sqrt{n}$ is $\rho(1-\rho)\textsf{Var}(m-\textsf{Pj}[m|\bm{X}])$ which is controlled by $\rho(1-\rho)\textsf{Var}(m)$. The procedure will not cause the ”shift problem” and the inflation of variance.
@@
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{2}]=O\big(n^{\frac{1}{2(r_{0}-1)}}\big)=o(n)$.

### v2

@@ Abstract
Covariate adaptive randomization (CAR) procedures are extensively used to reduce the likelihood of covariate imbalances occurring in clinical trials. In literatures, a lot of CAR procedures have been proposed so that the specified covariates are balanced well between treatments. However, the variance of the imbalance of the unspecified covariates may be inflated comparing to the one under the simple randomization. The variance inflation causes the usual test of treatment effects being not valid and adjusting the test being not an easy work. In this paper, we propose a new kind of covariate adaptive randomization procedures to balance covariates between two treatments with a ratio $\rho:(1-\rho)$. Under this kind of CAR procedures, the convergence rate of the imbalance of the specified covariates is $o(n^{1/2})$, and at the same time the asymptotic variance of the imbalance of any unspecified (observed or unobserved) covariates does not exceed the one under the simple randomization. The ”shift problem” found by Liu, Hu, and Ma (2025) will not appear under the new CAR procedures.
@@ 1 Introduction
Covariate adaptive randomization (CAR) procedures have often been implemented for balancing treatment allocation over influential covariates in clinical trials (McEntegart, , 2003; Rosenberger and Lachin, , 2015; Taves, , 2010; Lin, Zhu and Su, , 2015). The main purposes are to enhance the comparability of treatment groups and to increase the efficiency of the statistical inference. Many CAR procedures have been proposed to balance treatment assignments within strata and over covariate margins, provided that the covariates under consideration are discrete (categorical with two or more levels) (Pocock and Simon, , 1975; Antognini and Zagoraiou, , 2011; Hu and Hu, , 2012; Rosenberger and Lachin, , 2015). The theories on CAR randomization procedures are established progressively in the past decade(Hu and Hu, , 2012; Ma, Hu and Zhang, , 2015; Hu and Zhang, , 2020; Hu, Ye and Zhang, , 2023). For balancing continuous covariates, besides the method of discretization, a variety of methods have been proposed by minimization of prescribed imbalance measures (Hoehler, , 1987; Nishi and Takaichi, , 2004; Endo et al., , 2006; Stigsby and Taves, , 2010; Ciolino et al., , 2011; Lin and Su, , 2012; Ma and Hu, , 2013; Qin et al., , 2018; Jiang, Ma and Yin, , 2018; Li, Zhou and Hu, , 2019). Recently, Ma et al, (2024) unified a lot of CAR procedures by proposing a general CAR procedure for balancing specified covariate features $\bm{\phi}(\bm{X}_{i})$ between two treatments $1$ and $2$, where $\bm{\phi}(\bm{X}_{i})=(\phi_{1}(\bm{X}_{i}),\ldots,\phi_{q}(\bm{X}_{i}))$ can contain additional covariate features besides the covariate vector $\bm{X}_{i}=(X_{i,1},\ldots,X_{i,p})$ of a unit in clinical trals, such as quadratic and interaction terms. It is shown that the convergence rate of the imbalance vectors $\sum_{i=1}^{n}(T_{i}-1/2)\bm{\phi}(\bm{X}_{i})$ is $o_{P}(n^{\epsilon_{0}})$, $\epsilon_{0}<1/2$, where $n$ is the sample size, $\bm{X}_{i}$ is the covariate vector of the $i$-th unit, and $T_{i}=1$, 0 stands for the $i$-unit being allocated to treatment $1$ and $2$, respectively. Under some suitable conditions on the density of $\bm{\phi}(\bm{X}_{i})$, $\sum_{i=1}^{n}(T_{i}-1/2)\bm{\phi}(\bm{X}_{i})$ can attain the fastest convergence rate $O_{P}(1)$, and at the same time, for any unspecified covariate feature $m(\bm{X}_{i})$, the asymptotic normality of its normalized imbalance $\sum_{i=1}^{n}(T_{i}-1/2)m(\bm{X}_{i})/\sqrt{n}$ holds and the asymptotic variance $\sigma_{m}^{2}$ is nonzero when $m(\bm{X}_{i})$ is not a linear function of $\bm{\phi}(\bm{X}_{i})$. However, $\sigma_{m}^{2}$ may exceed $\textsf{E}[m^{2}(\bm{X}_{i})]/4$ that is the variance under the complete randomization in which each unit is allocated to treatments $1$ and $2$ with equal probability $1/2$. This issue is referred to as the variance inflation. The variance inflation under the procedure of Pocock and Simon, (1975) is also founded by Zhao et al., (2024). Because the asymptotic variance of the test of treatment effects is usually a function of $\sigma_{m}^{2}$ for some unspecified covariate feature $m(\bm{X}_{i})$, the variance inflation causes such test being not valid (Ma et al, , 2024; Zhang, , 2023). The closed form of $\sigma_{m}^{2}$ is also unknown so that estimating it and adjusting the test are not easy works. A lot of methods for the adjustment to obtain robust statistical inferences in the CAR clinical trials have being proposed(Bugni, Canay and Shaikh, , 2018; Ye and Shao, , 2020; Liu and Yang, , 2020; Ma, Tu and Liu, , 2024; Wang et al., , 2023). The purpose of this paper is to propose a new kind of CAR procedures so that the specified covariate features $\bm{\phi}(\bm{X}_{i})$ can be balanced well and at the same time the asymptotic variance $\sigma_{m}^{2}$ of the normalized imbalance of any unspecified covariate feature $m(\bm{X}_{i})$ is not inflated and has a closed form.
@@
From the proofs (Hu and Zhang, , 2020; Ma et al, , 2024; Zhang, , 2023), we can find that the convergence of $\sum_{i=1}^{n}(T_{i}-1/2)m(\bm{X}_{i})/\sqrt{n}$ and the finiteness of its asymptotic variance $\sigma_{m}^{2}$ are strongly based on the theory of Markov chains and the property that the allocation probabilities are symmetric about treatments so that, as a Markov chain, $\sum_{i=1}^{n}(T_{i}-1/2)\bm{\phi}(\bm{X}_{i})$ has a symmetric invariant distribution and thus $(T_{i}-1/2)m(\bm{X}_{i})$s have zero means under the invariance distribution for any $m(\bm{X}_{i})$. Recently, Liu, Hu and Ma, (2025) generalized Ma et al, (2024)’s procedure to balance covariances between two treatments with a ratio $\rho:(1-\rho)$ instead of half to half. The convergence rate of the imbalance vectors $\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})$ of the specified covariate feature $\bm{\phi}(\bm{X}_{i})$ is also shown to be $o_{P}(n^{\epsilon_{0}})$, $\epsilon_{0}<1/2$. However, by their theory, for an unspecified covariate feature $m(\bm{X})$, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})/n$ may converge to a nonzero constant $c_{m}$ due to the fact that $\sum_{i=1}^{n}(T_{i}-\rho)\bm{\phi}(\bm{X}_{i})$ is no longer a symmetric Markov chain when $\rho\neq 1/2$. This issue is refer to as ”shift problem” of covariate $m(\bm{X}_{i})$. The shift problem is a bad news because in testing hypotheses in the covaraite-adaptive randomized clinical trials, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})=o_{P}(n)$ is a basic condition (Ma, Hu and Zhang, , 2015; Ma et al., , 2020; Ma et al, , 2024; Zhu et al, , 2025; Liu and Hu, , 2023; Liu, Hu and Ma, , 2025), and sometimes the convergence rate $O_{P}(\sqrt{n})$ is needed (Ma, Tu and Liu, , 2024). We will show that our proposed CAR procedures are valid for all the case of $0<\rho<1$. Under the proposed CAR procedure, the ”shift problem” will no longer appear and the asymptotic variance of $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})/\sqrt{n}$ will be always controlled by $\rho(1-\rho)\textsf{E}[m^{2}(\bm{X}_{i})]$, the variance under the simple randomization in which each unit is randomized to treatments $1$ and $2$ respectively with $\rho$ and $1-\rho$, for any $m(\bm{X}_{i})$.
@@ 2.1 Framework
    $\ell_{n}=:\textsf{P}(T_{n}=1|\bm{X}_{n},...,\bm{X}_{1},T_{n-1},...,T_{1})=\ell\left(\frac{\langle\bm{\Lambda}_{n-1},\bm{\phi}(\bm{X}_{n})\rangle}{(n-1)^{\gamma}}\right),$ (2.1)
@@
The function $\ell(x)$ in (2.1) is called the allocation function. A proposed choice of $\ell$ is
where $\Phi$ is the standard normal distribution, $x\wedge y=\min\{x,y\}$ and $x\vee y=\max\{x,y\}$. Another simple choice of $\ell$ is
@@
According to the theoretical results given in Theorem 2.2 and Remark 2.2, the parameter $\gamma$ controls the size of the imbalance measure $Imb_{n}$. The smaller the parameter $\gamma$ , the smaller is $Imb_{n}$, but possibly the larger is the imbalance $\|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\|^{2}$ of the unspecified feature $Z_{i}$ that is not a linear fuction of $\bm{\phi}(\bm{X}_{i})$. The value of $\gamma$ is suggested to be in $[0.5,1)$ so that the asymptotic expectation of $\|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\|^{2}/n$ does not exceed the one under the simple randomization for any $Z_{i}$s.
@@ 2.2 Theoretical Properties
The sequence $\{(\bm{X}_{i},Z_{i},W_{i})\}$ is a sequence of independent and identically distributed as a random vector $(\bm{X},W,Z)$..
@@
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{2}]=O\big(n^{\gamma}+n^{\frac{1}{2}+\frac{1}{2(r_{0}-1)}}\big)=o(n).$ (2.5)
If $\textsf{E}[|Z|]<\infty$ and $\gamma>\frac{1}{4}+\frac{1}{4(r_{0}-1)}$, then
$\textsf{E}|(\ell_{n}-\rho)Z_{n}|\to 0,$ (2.6)
$\frac{\textsf{E}\left|\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}\right|}{n}\to 0.$ (2.7)
Moreover, if $r_{0}\geq 4$, $\gamma>\frac{1}{2}+\frac{1}{4(r_{0}-1)}$, $\textsf{E}Z^{2}<\infty$, $\textsf{E}W^{2}<\infty$ and $\textsf{E}[W]=0$, then
@@
(i) No shift problem:
Liu, Hu and Ma, (2025) proposes a CAR procedure for balancing covariate features $\bm{\phi}(\bm{X}_{i})$ between treatments with an unequal ratio $\rho:(1-\rho)$. When the additional covariate $m(\bm{X}_{i})$ is not the linear function of $\bm{\phi}(\bm{X}_{i})$, it may happen that $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})\overset{P}{\to}c\neq 0$. This issue is referred as ”shift problem”. The simulation results presented in the Supplementary Material of Liu, Hu and Ma, (2025) show that the shift problem on the additional covariates is common when balancing continuous covariates by using their CAR procedure. Now, (2.7) shows that the shift problem no longer exists under our proposed CAR procedure.
(ii) Closed form of variance:
$\langle\bm{\phi}(\bm{X}),\bm{x}_{0}\rangle$ is the orthogonal projection of $Z$ on $\bm{\phi}(\bm{X})$, we denote it by $\textsf{Pj}[Z|\bm{\phi}(\bm{X})]$. Thus, $\vec{\sigma}_{Z}^{2}=\textsf{E}\big(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})]\big)^{2}$. This closed form makes $\vec{\sigma}_{Z}^{2}$ can be estimated basing on the observations of $Z$ and $\bm{\phi}(\bm{X})$.
It is easily seen that $\textsf{E}\big(Z-\textsf{Pj}[Z|\bm{\phi}(\bm{X})]\big)^{2}=\textsf{E}Z^{2}-\textsf{E}\big(\textsf{Pj}[Z|\bm{\phi}(\bm{X})]\big)^{2}\leq\textsf{E}Z^{2}$. Thus, $\vec{\sigma}_{Z}^{2}=0$ if and only if $Z\in Span\{\bm{\phi}(\bm{X})\}=:\{\langle\bm{x},\bm{\phi}(\bm{X})\rangle:\bm{\in}\mathbb{R}^{q}\}$, and $\vec{\sigma}_{Z}^{2}=(1-\rho)\rho\textsf{E}[Z^{2}]$ if and only if $\textsf{E}[\bm{\phi}(\bm{X})Z]=\bm{0}$, i.e., $Z\perp\bm{\phi}(\bm{X})$. As a result, $\vec{\sigma}_{Z}^{2}=\textsf{Var}(Z-\textsf{E}[Z|\bm{\phi}(\bm{X})])+\vec{\sigma}_{\textsf{E}[Z|\bm{\phi}(\bm{X})]}^{2}$. Further, if $1\in Span\{\bm{\phi}(\bm{X})\}$, then $\vec{\sigma}_{Z}^{2}=\vec{\sigma}_{Z-\textsf{E}Z}\leq\textsf{Var}(Z)$.
(iii)No variance inflation:
The simple randomization is randomizing each patient to treatment $1$ with the same probability $\rho$. For the simple randomization, $\vec{\sigma}_{Z}^{2}=\rho(1-\rho)\textsf{E}Z^{2}$. So, the proposed CAR procedure always has smaller asymptotic variance than the simple randomization for all observed or unobserved feature $Z_{i}$ with $\textsf{Pj}[Z|\bm{\phi}(\bm{X})]\neq 0$. The variance inflation is no longer a problem under our proposed CAR procedure.
(iv) Independence of exogenous variables:
The normalized imbalance $\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}/\sqrt{n}$ of $Z_{i}$s is an endogenous variable of the design, and the normalized sum $\sum_{i=1}^{n}W_{i}/\sqrt{n}$ of $W_{i}$s can be regraded as an exogenous variable. (2.8) means that they are asymptotic independent regardless $W_{i}$ and $Z_{i}$ themselves are independent or not.
When $\bm{\phi}(\bm{X})$ has finite moments of all orders. The convergence rate of the imbalance $\bm{\Lambda}_{n}$ can be improved a lot.
@@
Suppose Assumption 2.1, and that 2.2 is satisfied for all $r_{0}>2$.
(2.6) and(2.7) hold if $\textsf{E}[|Z|]<\infty$, and
$\big(\textsf{E}\|\bm{\Lambda}_{n}\|^{r}\big)^{1/r}=O(n^{\gamma/2})\;\text{ for all }r>0.$ (2.9)
If $\textsf{E}Z^{2}<\infty$, then
@@
for all $r>0$, where $\widetilde{Z}_{i}=Z_{i}-\textsf{Pj}[Z_{i}|\bm{\phi}(\bm{X}_{i})]$.
If $\textsf{E}Z^{2}<\infty$, $\textsf{E}W^{2}<\infty$, $\textsf{E}W=0$ and $1/2<\gamma<1$, then (2.8) holds,
@@
If $\ell(x)$ is thrice differentiable at $x=0$, and $\ell^{\prime\prime}(0)=0$ or $\textsf{E}[(\bm{\phi}(\bm{X}))^{\prime}\bm{\phi}(\bm{X})\widetilde{Z}]=\bm{0}$, then $n^{1-\gamma}$ in ((ii)) can be repalced by $n^{1-3\gamma/2}$, and (2.8), (2.11), (2.12) remain true when $1/3<\gamma<1$.
@@
where $\lambda>0$ is a tuning parameter. In (2.13), $\Phi(-x+u_{\rho/2})$, $\Phi(-x+u_{(\rho+1)/2})$ can be replaced by $\alpha\rho$, $1-(1-\alpha)\rho$, respectively, $0<\alpha<1$.
@@
where $w_{o}$, $w_{m,l}$, $w_{s}$ are the corresponding nonnegative weights and at least one of them is no zero.
@@
$=$ $w_{0}D_{n-1}+\sum_{l=1}^{p}w_{m,l}D_{n-1}(l;k_{l}^{\ast})+w_{s}D_{n-1}(k_{1}^{\ast},\ldots,k_{p}^{\ast}).$
Then, the allocation probability of the $n$th-unit is
@@
Thus, when $w_{s}\neq 0$, $D_{n}^{2}(k_{1},\ldots,k_{p})=O_{P}(n^{\gamma/2})$ and so for any $m(\bm{X})$, $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})=O_{P}(n^{\gamma/2})$. When $w_{s}=0$, we will have
@@
When $\rho=1/2$ and $\gamma=0$, our CRA procedure is similar to that of Hu and Hu, (2012). The theory of this kind of CAR procedures can be found in Hu and Zhang, (2020) and Zhang, (2023). It is shown that (2.17) still holds, but $\vec{\sigma}_{m}^{2}$ has no closed from and may exceed $\textsf{E}[m^{2}(\bm{X})]/4$ unless $\textsf{E}[m(\bm{X})|\bm{\phi}(\bm{X})]\in Span\{\bm{\phi}(\bm{X})\}$. In particular, when $w_{s}=w_{0}=0$, the procedure is Pocock and Simon, (1975)’s marginal procedure. Zhao et al., (2024) demonstrated via examples that along some directions, there is an increased variance of within-stratum imbalances under Pocock and Simon, (1975)’s procedure compared to the simple randomization.
@@
when $X_{1},\ldots,X_{p}$ have finite moments of each order. At the same time, for any nonlinear function $m=m(X_{1},\ldots,X_{p})$ with finite variance, then the asymptotic variance of $\sum_{i=1}^{n}(T_{i}-\rho)m(\bm{X}_{i})/\sqrt{n}$ is $\rho(1-\rho)\textsf{Var}(m-\textsf{Pj}[m|\bm{X}])$ which is controlled by $\rho(1-\rho)\textsf{Var}(m)$. The procedure will not cause the ”shift problem” and the variance inflation.
@@
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{2}]=O\big(n^{\frac{1}{r_{0}-1}}\big)$, and $O\big(n^{\frac{8}{(r_{0}+1)^{2}}}\big)$ when $r_{0}\geq 3$.